OpenStudy (anonymous):

Calc III- Please help! A lamina occupies the part of the disk x^2+y^2<=1 in the first quadrant. Find its center of mass if the density at any point is proportional to the square of the distance from the origin.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@jim_thompson5910 @amistre64

OpenStudy (anonymous):

@iambatman @aaronq

OpenStudy (anonymous):

\[x ^{2}+y ^{2}\le1\]

OpenStudy (perl):

$$ \Large \iint_D \rho (x,y) dA $$

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

but im having troubel setting up the density function

OpenStudy (perl):

it would be easier to convert this problem to polar coordinates

OpenStudy (anonymous):

ok so r=1

OpenStudy (anonymous):

or r<=1

OpenStudy (perl):



OpenStudy (perl):

$$ \Large \rho (x,y) = k (\sqrt{ x^2 + y^2 })^2 $$

OpenStudy (anonymous):

ahh i see

OpenStudy (perl):

The region is described as the set of all \( (r, \theta ) \) such that \( 0\leq r\leq 1 \) and \( 0 \leq \theta \leq \pi/2 \).

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{2\pi}kr ^{3}d thetadr\]

OpenStudy (anonymous):

wait the limit is pi/2 not 2pi

OpenStudy (perl):

right, that will give you the mass . and then you need moment with respect to x axis and moment with respect to y axis .

OpenStudy (anonymous):

its just that times x and y right?

OpenStudy (dan815):